Lemon drops and jelly beans are mixed to make a 100 pound mix. If the lemon drops cost $1.90 per pound and jelly beans cost $1.20 per pound, how many pounds of lemon drops are needed to make the mix cost $1.48 per pound?
Answer by richwmiller(17219)

l=lemon drops
j=jelly beans
l+j=100
l*1.90+j*1.20=100*148
j=60 and l=40